# SBP-FDEC: Summation-by-Parts Finite Difference Exterior Calculus

### v2

@@ 1 Introduction
The theory of Finite Element Exterior Calculus (FEEC) by Arnold, Falk, and Winther [feec1, feec2], Arnold [feec3], with foundational contributions by Monk [peter_maxwell], is a state-of-the-art finite element framework that uses the structure of the de Rham cohomology to preserve topological structures, such as the cohomology groups of the region, or manifold, on which the Partial Differential Equations (PDEs) in question are posed. The primary example used in these foundational papers is the Hodge-Laplacian, but another major application of this framework are problems where an exactly divergence- or curl-free solution is needed. For instance, FEEC has been used to satisfy the divergence-free condition of the magnetic field when discretizing the time-dependent Maxwell’s equations (see, for instance, [peter_maxwell] and [feec3]) and Vlasov–Maxwell equations (see, for instance [gempic, Campos-Pinto2022Variational]). The theory has also been applied to other PDEs that need to satisfy divergence constraints, such as the incompressible Navier–Stokes equations. See, for example, Palha et al. [palha_navier_stokes] or Carlier et al. [carlier_navier_stokes]. In FEEC, compatible ansatz spaces and projections are used to discretize the PDEs. For that purpose, two (for 1D), three (for 2D), or four (in 3D) discrete spaces and conforming projections onto these spaces are necessary to discretize the de Rham cochain complex. The FEEC framework has also recently been extended to broken spaces, such that only local mass matrices are needed [gucclu2023broken].
@@
In the literature, a variety of different ansatz spaces and projections have been proposed to discretize the de Rham cochain complex, most of which are polynomial or piecewise polynomial in nature. Classical examples include Nédélec spaces of the first and second kind [nedelec1, nedelec2], Whitney forms [whitney] or Raviart–Thomas elements [raviart_thomas]. These predate the formulation of the FEEC framework. More recently, spline elements using a tensor-product ansatz on hexahedral spaces have also been used, for example, in Buffa et al. [isogeometric_buffa] and Evans et al. [evans_isogeometric].
@@
These approaches differ not only in the choice of finite element spaces, but also in the projection operators they employ, which leads to variations in the associated degrees of freedoms and basis functions of each approach. For instance, the approach of Gerritsma [gerritsma_histopolation] and Gerritsma et al. [gerritsma_preprint], which motivated the present work, uses nodal and integral degrees of freedom on a one-dimensional subgrid per element. Using tensor products, this construction extends to nodal, line-integral, area-integral, and volume-integral degrees of freedom on a multi-dimensional level. Other approaches, as in Nédélec [nedelec1, nedelec2], use higher moment degrees of freedom instead of a subgrid approach.
@@
There is a rich literature of discrete operators on finite-difference-like degrees of freedom using discrete derivative operators to mimic the de Rham complex discretely to achieve similar structure preserving properties as FEEC. These schemes are also known as mimetic finite-difference methods. One survey paper by Lipnikov et al. [lipnikov] gives a comprehensive overview of these methods. An early example of such schemes is given in the book of Samarskii [samarskii_translation]. Another important work is the method of Yee [yee] for Maxwell’s equations, which is the foundation for the class of finite-difference time-domain methods. It uses a staggered space-time grid for the different fields, which automatically preserves the divergence condition of the magnetic field. Another approach of Hyman and Shashkov in [hyman_discrete_adjoints, hyman_discrete_operators] and [hyman_boundaries] uses the derivation of discrete analogues for the grad, curl, and div operators and their adjoints. In addition to the natural operators, which are these discrete analogues, arising in this context, mimetic discretizations rely on a reconstruction operator to compute point values of the different fields involved. For high-order mimetic finite differences, these are mostly interpolation operators based on the degrees of freedoms in the neighboring cells, as for example in high-order finite volume reconstructions. In [Kormann2024A-Dual-Grid] a mimetic method on staggered grids is developed for the Vlasov–Maxwell equations, where the reconstructions are defined by a sliding stencil of Lagrange interpolation or histopolation using the neighboring cells.
@@
A different class of finite-difference methods, called Summation-by-Parts Finite Differences (SBP-FD), introduced Kreiss and Scherer [KREISS_1974] with key theoretical contributions from Strand [strand_sbp], were developed to re-create energy estimates of classic Finite-Element Methods (FEM). SBP operators use a discrete norm (mass) matrix together with a derivative matrix that operate on nodal degrees of freedom. These operators are constructed such that they mimic integration by parts discretely, enabling the derivation of discrete energy estimated, see Olsson in [olsson_sbp_1, olsson_sbp_2], and entropy estimates, see for instance Crean et al. [sbp_entropy_crean] or Carpenter et al. [sbp_entropy_carpenter]. A defining feature of these methods is the equivalence of their strong and weak formulations, which is central to obtaining stability results.
@@
These schemes can be extended to a multi-element formulation with the Simultaneous Approxmation Term (SAT) approach, treating interface and boundary conditions weakly. For further information, see, for instance, the review papers [review_sbp_sat, fernandez2014review]. SBP operators for the first and second derivative have also been developed for non-polynomial approximation spaces, for instance for trigonometric functions or radial basis functions, see Glaubitz et al. [generalized_function_spaces_1, generalized_function_spaces_2]. The SBP framework has also influenced methods beyond finite-difference schemes. One example of this is the Discontinuous Galerkin Spectral Element Method (DGSEM) using a Legendre–Gauss–Lobatto (LGL) grid, see Kopriva [dgsem], where the mass and derivative matrices form an SBP operator, Gassner [gassner_skew_burgers].
@@
To address this issue, we exploit important connections between DGSEM-LGL and SBP-FD, and borrow important ingredients from the work of Gerritsma [gerritsma_preprint] and Gerritsma et al. [gerritsma_histopolation], who developed FEEC operators using LGL spectral element operators. In Gerritsma et al.’s approach, a subgrid structure is imposed on each LGL element, accompanied by two compatible polynomial spaces and two projection operators in 1D: (i) the polynomial interpolation space and (ii) the polynomial histopolation space.
@@
The remainder of the paper is organized as follows. In section 2, we revisit and generalize the ansatz of Gerritsma [gerritsma_histopolation] to the SBP-FD framework. We include extensions to multiple elements and derivations of discrete 2D and 3D derivative operators based on existing SBP-FD operators. In section 3, we apply this theory to the 2D transverse electric homogeneous Maxwell’s equation to demonstrate the capability of the SBP-FDEC to exactly preserve the divergence-free electric field. In section 4, we derive the weak and strong form of the SBP-FDEC method for the Maxwell’s equations on a periodic domain and prove the discrete divergence-free property and semi-discrete energy conservation, which can then be extended to fully discrete energy conservation using appropriate established time-integration methods. In section 5, we numerically verify our method on a periodic test case, showing the experimental order of convergence fitting to the established SBP-FD operators, demonstrating that the expected convergence rate is obtained as well as that the scheme is discretely divergence-free and energy preserving. The last section concludes the paper.
@@ 2 Mimetic Finite Difference Operators
To obtain such a discrete complex, we will generalize the ansatz proposed by Gerritsma [gerritsma_histopolation], who derives two compatible polynomial spaces using nodal and integral degrees of freedom on a given grid. Instead of using polynomial spaces, however, we assume an unknown general Lagrange basis that we only know through discrete nodal function values and the corresponding discrete nodal derivative values. This assumption allows us to connect nodal function values to a new set of discrete integral degrees of freedom and further derive corresponding discrete representations and matrix operators.
@@
We call these functions edge functions or histopolation functions associated with the Lagrange functions $\{l_{i}\}_{i=0}^{N}$, using the terminology of Gerritsma [gerritsma_histopolation]. It is straightforward to check that the histopolation property
@@
holds. This property is analogous to the Lagrange property on the grid nodes but instead applies to the integrals over the sub-intervals defined by these nodes. It follows that $\{h_{i}\}_{i=0}^{N}$ are linearly independent functions and the span of these functions is a space of dimension $N$. Since the $h_{i}$ are sums of derivatives of the Lagrange functions, we know that $h_{i}\in U_{h}$ for all $i\in\{1,\dots,N\}$. On the other hand, the dimension of $U_{l}$ is $N+1$ and $U_{l}$ contains all constant functions. Hence, $\text{dim}\left(U_{h}\right)=\text{dim}\left(\frac{\partial{}}{\partial{x}}\left(U_{l}\right)\right)=N$, which shows that $U_{h}$ is the span of the histopolation functions. In addition, we know that all constant functions are contained in $U_{h}$ since the linear functions are contained in $U_{l}$. This is equivalent to
$\sum_{i=1}^{N}\left(x_{i+1}-x_{i}\right)h_{i}(x)=1,\qquad\forall x\in[a,b].\\$ (6)
Given a function $g\in U_{h}$, the histopolation property (5) of the edge functions implies that the functions $h_{i}$ form a complete basis for $U_{h}$, i.e.
@@
Let the interpolation operator $p^{1}$ and the histopolation operator $p^{0}$ be defined as
@@
$\frac{\partial{}}{\partial{x}}p_{0}(f)(x_{i})$ $=\sum_{k=0}^{N}D_{i,k}f_{k}$ $=$ $\left(\underline{D}\textbf{f}\right)_{i}$
$=-\sum_{j=1}^{N}\left(\left(f_{j}-f_{j-1}\right)\sum_{k=0}^{j-1}D_{i,k}\right)$ $=$ $\left(\underline{V}\,\underline{\Delta}\textbf{f}\right)_{i}=p_{1}\left(\frac{\partial{}}{\partial{x}}f\right)(x_{i})$
@@
$\underline{\Delta}=\begin{pmatrix}-1&1&0&0&\dots\\ 0&-1&1&0&\dots\\ 0&0&-1&1&\dots\\ \vdots&\vdots&\vdots&\vdots&\ddots\\ \end{pmatrix}$ (16)
@@ 2.2 Derivative Operators
The main idea of what follows is motivated by the work of Gerritsma [gerritsma_histopolation], and Gerritsma et al. [gerritsma_preprint] which we aim to combine with our findings on the discrete relationships of $\underline{D}$, $\underline{V}$, and $\underline{\Delta}$ for general Lagrange and histopolation basis functions, which do not have to be polynomials. In particular, we are interested in high-order SBP-FD operators, for which we know the basis function values and their derivatives at the grid points. In the following, vector-valued derivative operators will be denoted by bold fonts.
@@
$\frac{\partial{}}{\partial{x}}\phi(x,y,z)=\sum_{i=1}^{N}\sum_{j=0}^{N}\sum_{k=0}^{N}(\phi_{i,j,k}-\phi_{i-1,j,k})h_{i}(x)l_{j}(y)l_{k}(z).$ (22)
@@
As described in [gerritsma_histopolation], the histopolation derivative (22) can be used directly to define a discrete grad operator from the space $\left(U_{l}\otimes U_{l}\otimes U_{l}\right)^{T}$ to the space $(U_{h}\otimes U_{l}\otimes U_{l},U_{l}\otimes U_{h}\otimes U_{l},U_{l}\otimes U_{l}\otimes U_{h})$. The approximation spaces for the curl and div operators are adjusted accordingly with tensor product combinations of Lagrange and histopolation functions, where the general rule is that the directions in which derivatives are applied are represented by $U_{l}$ and other directions are represented by $U_{h}$. For a more detailed explanation, see [gerritsma_preprint].
As mentioned, our goal is to obtain purely discrete versions of these approximations at grid nodes for our mimetic FD approximations. It is straight forward to use the relationships in section 2.1 to obtain analogous expressions for grad, curl, and div that are purely nodal-based. We do note however, that the underlying degrees of freedom are in general not only nodal, but also consist of line, area and volume integral degrees of freedom of the cell complex induced by the tensor product grid. The derivative matrix however gives us a way of computing nodal values from the integral degrees of freedom.
@@
$\frac{\partial{}}{\partial{x}}\phi(x_{i},y_{j},z_{k})=\sum_{s=1}^{N}V_{i,s}\left(\phi_{s,j,k}-\phi_{s-1,j,k}\right)=\sum_{s=0}^{N}D_{i,j}\phi_{s,j,k}.$ (23)
@@
In the following, we present discrete versions of all operators described by Gerritsma [gerritsma_histopolation], which can be computed using arbitrary Lagrange and histopolation basis functions, for which we only need to know the function values and their derivatives at the grid nodes. On the other hand, as long as the values of the basis functions at the grid nodes are known, we can build the operators without having access to explicit basis functions. For instance, we can use high-order SBP-FD operators.
@@ 2.2.2 Curl Operator
We aim to define the curl operator in a space compatible with the gradient operator described above, so that we preserve the well-known identity that the curl of the gradient of a scalar field is zero. We choose $\tilde{\textbf{u}}=(\tilde{u}^{x},\tilde{u}^{y},\tilde{u}^{z})$ from the same space as $\textbf{grad}(\phi)$ defined above. I.e., we have $\tilde{u}^{x}$ is from the tensor product space $U_{h}\otimes U_{l}\otimes U_{l}$, $\tilde{u}^{y}$ from $U_{l}\otimes U_{h}\otimes U_{l}$, and $\tilde{u}^{z}$ from $U_{l}\otimes U_{l}\otimes U_{h}$. Then the nodal values of $\textbf{curl}(\tilde{\textbf{u}})=\textbf{w}=(w^{x},w^{y},w^{z})^{T}$ are given by
$\begin{split}w^{x}(x_{i},y_{j},z_{k})&=\sum_{s=1}^{N}\sum_{t=1}^{N}(\tilde{u}_{i,s,t}^{z}-\tilde{u}_{i,s-1,t}^{z}-\tilde{u}_{i,s,t}^{y}+\tilde{u}_{i,s,t-1}^{y})V_{j,s}V_{k,t}\\ &=\sum_{s=1}^{N}\sum_{t=1}^{N}\tilde{u}_{i,s,t}^{z}D_{j,s}V_{k,t}-\tilde{u}_{i,s,t}^{y}V_{j,s}D_{k,t},\end{split}$ (27)
$\begin{split}w^{y}(x_{i},y_{j},z_{k})&=\sum_{s=1}^{N}\sum_{t=1}^{N}(\tilde{u}_{s,j,t}^{x}-\tilde{u}_{s,j,t-1}^{x}-\tilde{u}_{s,j,t}^{z}+\tilde{u}_{s-1,j,t}^{z}V_{i,s}V_{k,t}\\ &=\sum_{s=1}^{N}\sum_{t=1}^{N}\tilde{u}_{s,j,t}^{x}V_{i,s}D_{k,t}-\tilde{u}_{s,j,t}^{z}D_{i,s}V_{k,t},\end{split}$ (28)
$\begin{split}w^{z}(x_{i},y_{j},z_{k})&=\sum_{s=1}^{N}\sum_{t=1}^{N}(\tilde{u}_{s,t,k}^{y}-\tilde{u}_{s-1,t,k}^{y}-\tilde{u}_{s,t,k}^{x}+\tilde{u}_{s,t-1,k}^{x})V_{i,s}V_{j,t}\\ &=\sum_{s=1}^{N}\sum_{t=1}^{N}\tilde{u}_{s,t,k}^{y}D_{i,s}V_{j,t}-\tilde{u}_{s,t,k}^{x}V_{i,s}D_{j,t}.\end{split}$ (29)
@@
$\begin{split}&u_{i,j}^{k}-u_{i,j-1}^{k}-u_{i,k}^{j}+u_{i,k-1}^{j}\\ =&\phi_{i,j,k}-\phi_{i,j,k-1}-\phi_{i,j-1,k}+\phi_{i,j-1,k-1}\\ -&\phi_{i,j,k}+\phi_{i,j-1,k}+\phi_{i,j,k-1}-\phi_{i,j-1,k-1}\\ =&0,\end{split}$ (30)

### v3

@@ 1 Introduction
The theory of Finite Element Exterior Calculus (FEEC) by Arnold, Falk, and Winther [2, 3], Arnold [4], with foundational contributions by Monk [32], is a state-of-the-art finite element framework that uses the structure of the de Rham cohomology to preserve topological structures, such as the cohomology groups of the region, or manifold, on which the Partial Differential Equations (PDEs) in question are posed. The primary example used in these foundational papers is the Hodge-Laplacian, but another major application of this framework are problems where an exactly divergence- or curl-free solution is needed. For instance, FEEC has been used to satisfy the divergence-free condition of the magnetic field when discretizing the time-dependent Maxwell’s equations (see, for instance, [32] and [4]) and Vlasov–Maxwell equations (see, for instance [26, 7]). The theory has also been applied to other PDEs that need to satisfy divergence constraints, such as the incompressible Navier–Stokes equations. See, for example, Palha et al. [38] or Carlier et al. [8]. In FEEC, compatible ansatz spaces and projections are used to discretize the PDEs. For that purpose, two (for 1D), three (for 2D), or four (in 3D) discrete spaces and conforming projections onto these spaces are necessary to discretize the de Rham cochain complex. The FEEC framework has also recently been extended to broken spaces, such that only local mass matrices are needed [18].
@@
In the literature, a variety of different ansatz spaces and projections have been proposed to discretize the de Rham cochain complex, most of which are polynomial or piecewise polynomial in nature. Classical examples include Nédélec spaces of the first and second kind [34, 35], Whitney forms [46] or Raviart–Thomas elements [40]. These predate the formulation of the FEEC framework. More recently, spline elements using a tensor-product ansatz on hexahedral spaces have also been used, for example, in Buffa et al. [6] and Evans et al. [13].
@@
These approaches differ not only in the choice of finite element spaces, but also in the projection operators they employ, which leads to variations in the associated degrees of freedoms and basis functions of each approach. For instance, the approach of Gerritsma [15] and Gerritsma et al. [27], which motivated the present work, uses nodal and integral degrees of freedom on a one-dimensional subgrid per element. Using tensor products, this construction extends to nodal, line-integral, area-integral, and volume-integral degrees of freedom on a multi-dimensional level. Other approaches, as in Nédélec [34, 35], use higher moment degrees of freedom instead of a subgrid approach.
@@
There is a rich literature of discrete operators on finite-difference-like degrees of freedom using discrete derivative operators to mimic the de Rham complex discretely to achieve similar structure preserving properties as FEEC. These schemes are also known as mimetic finite-difference methods. One survey paper by Lipnikov et al. [29] gives a comprehensive overview of these methods. An early example of such schemes is given in the book of Samarskii [42]. Another important work is the method of Yee [47] for Maxwell’s equations, which is the foundation for the class of finite-difference time-domain methods. It uses a staggered space-time grid for the different fields, which automatically preserves the divergence condition of the magnetic field. Another approach of Hyman and Shashkov in [22, 20] and [21] uses the derivation of discrete analogues for the grad, curl, and div operators and their adjoints. In addition to the natural operators, which are these discrete analogues, arising in this context, mimetic discretizations rely on a reconstruction operator to compute point values of the different fields involved. For high-order mimetic finite differences, these are mostly interpolation operators based on the degrees of freedoms in the neighboring cells, as for example in high-order finite volume reconstructions. In [25] a mimetic method on staggered grids is developed for the Vlasov–Maxwell equations, where the reconstructions are defined by a sliding stencil of Lagrange interpolation or histopolation using the neighboring cells.
@@
A different class of finite-difference methods, called Summation-by-Parts Finite Differences (SBP-FD), introduced Kreiss and Scherer [28] with key theoretical contributions from Strand [44], were developed to re-create energy estimates of classic Finite-Element Methods (FEM). SBP operators use a discrete norm (mass) matrix together with a derivative matrix that operate on nodal degrees of freedom. These operators are constructed such that they mimic integration by parts discretely, enabling the derivation of discrete energy estimates, see Olsson in [36, 37], and entropy estimates, see for instance Crean et al. [10] or Carpenter et al. [9]. A defining feature of these methods is the equivalence of their strong and weak formulations, which is central to obtaining stability results.
@@
These schemes can be extended to a multi-element formulation with the Simultaneous Approxmation Term (SAT) approach, treating interface and boundary conditions weakly. For further information, see, for instance, the review papers [45, 11]. SBP operators for the first and second derivative have also been developed for non-polynomial approximation spaces, for instance for trigonometric functions or radial basis functions, see Glaubitz et al. [17, 16]. The SBP framework has also influenced methods beyond finite-difference schemes. One example of this is the Discontinuous Galerkin Spectral Element Method (DGSEM) using a Legendre–Gauss–Lobatto (LGL) grid, see Black [5] and Kopriva [24], where the mass and derivative matrices form an SBP operator, see Gassner [14].
@@
To address this issue, we exploit important connections between DGSEM-LGL and SBP-FD, and borrow important ingredients from the work of Gerritsma [27] and Gerritsma et al. [15], who developed FEEC operators using LGL spectral element operators. In Gerritsma et al.’s approach, a subgrid structure is imposed on each LGL element, accompanied by two compatible polynomial spaces and two projection operators in 1D: (i) the polynomial interpolation space and (ii) the polynomial histopolation space.
@@
The remainder of the paper is organized as follows. In section 2, we revisit and generalize the ansatz of Gerritsma [15] to the SBP-FD framework. We include extensions to multiple elements and derivations of discrete 2D and 3D derivative operators based on existing SBP-FD operators. We further show that the cohomology spaces of the de Rham complexes for non-periodic rectangular domains without imposed boundary conditions are preserved by the discrete complexes. In section 3, we apply this theory to the 2D transverse electric homogeneous Maxwell’s equation to demonstrate the capability of the SBP-FDEC to exactly preserve the divergence-free electric field. In section 4, we derive the weak and strong form of the SBP-FDEC method for the Maxwell’s equations on a periodic domain and prove the discrete divergence-free property and semi-discrete energy conservation, which can then be extended to fully discrete energy conservation using appropriate established time-integration methods. In section 5, we numerically verify our method on a periodic test case, showing the experimental order of convergence fitting to the established SBP-FD operators, demonstrating that the expected convergence rate is obtained as well as that the scheme is discretely divergence-free and energy preserving. The last section concludes the paper.
@@ 2 Mimetic Finite Difference Operators
To obtain such a discrete complex, we will generalize the ansatz proposed by Gerritsma [15], who derives two compatible polynomial spaces using nodal and integral degrees of freedom on a given grid. Instead of using polynomial spaces, however, we assume an unknown general Lagrange basis that we only know through discrete nodal function values and the corresponding discrete nodal derivative values. This assumption allows us to connect nodal function values to a new set of discrete integral degrees of freedom and further derive corresponding discrete representations and matrix operators.
@@
We call these functions edge functions or histopolation functions associated with the Lagrange functions $\{l_{i}\}_{i=0}^{N}$, using the terminology of Gerritsma [15]. It is straightforward to check that the histopolation property
@@
holds. This property is analogous to the Lagrange property on the grid nodes but instead applies to the integrals over the sub-intervals defined by these nodes. It follows that $\{h_{i}\}_{i=1}^{N}$ are linearly independent functions and the span of these functions is a space of dimension $N$. Since the $h_{i}$ are sums of derivatives of the Lagrange functions, we know that $h_{i}\in U_{h}$ for all $i\in\{1,\dots,N\}$. On the other hand, the dimension of $U_{l}$ is $N+1$ and $U_{l}$ contains all constant functions. Hence, $\text{dim}\left(U_{h}\right)=\text{dim}\left(\frac{\partial{}}{\partial{x}}\left(U_{l}\right)\right)=N$, which shows that $U_{h}$ is the span of the histopolation functions. Given a function $g\in U_{h}$, the histopolation property (5) of the edge functions implies that the functions $h_{i}$ form a complete basis for $U_{h}$, i.e.
@@
In addition, we know that all constant functions are contained in $U_{h}$ since the linear functions are contained in $U_{l}$. For $g(x)\equiv 1$ we have due to (6)
$g(x)=\sum_{i=1}^{N}\left(\int_{x_{i-1}}^{x_{i}}1\text{ d}s\right)h_{i}(x)=\sum_{i=1}^{N}\left(x_{i+1}-x_{i}\right)h_{i}(x)=1,\qquad\forall x\in[a,b].\\$ (7)
Let the interpolation operator $p^{0}$ and the histopolation operator $p^{1}$ be defined as
@@
$\begin{split}&\frac{\partial{}}{\partial{x}}p_{0}(f)(x_{i})=\sum_{k=0}^{N}D_{i,k}f_{k}=\left(\underline{D}\textbf{f}\right)_{i}\\ =&-\sum_{j=1}^{N}\left(\left(f_{j}-f_{j-1}\right)\sum_{k=0}^{j-1}D_{i,k}\right)\\ =&\left(\underline{V}\,\underline{\Delta}\textbf{f}\right)_{i}=p_{1}\left(\frac{\partial{}}{\partial{x}}f\right)(x_{i})\end{split}$ (16)
@@
$\underline{\Delta}=\begin{pmatrix}-1&1&0&0&\dots\\ 0&-1&1&0&\dots\\ 0&0&-1&1&\dots\\ \vdots&\vdots&\vdots&\vdots&\ddots\\ \end{pmatrix}$. (17)
@@ 2.2 Derivative Operators
The main idea of what follows is motivated by the work of Gerritsma [15], and Gerritsma et al. [27] which we aim to combine with our findings on the discrete relationships of $\underline{D}$, $\underline{V}$, and $\underline{\Delta}$ for general Lagrange and histopolation basis functions, which do not have to be polynomials. In particular, we are interested in high-order SBP-FD operators, for which we know the basis function values and their derivatives at the grid points. In the following, vector-valued derivative operators will be denoted by bold fonts.
@@
$\frac{\partial{\phi}}{\partial{x}}(x,y,z)=\sum_{i=1}^{N}\sum_{j=0}^{N}\sum_{k=0}^{N}(\phi_{i,j,k}-\phi_{i-1,j,k})h_{i}(x)l_{j}(y)l_{k}(z).$ (23)
@@
As described in [15], the histopolation derivative (23) can be used directly to define a discrete grad operator from the space $\left(U_{l}\otimes U_{l}\otimes U_{l}\right)^{T}$ to the space $(U_{h}\otimes U_{l}\otimes U_{l},U_{l}\otimes U_{h}\otimes U_{l},U_{l}\otimes U_{l}\otimes U_{h})$. The approximation spaces for the curl and div operators are adjusted accordingly with tensor product combinations of Lagrange and histopolation functions, where the general rule is that the directions in which derivatives are applied are represented by $U_{l}$ and other directions are represented by $U_{h}$. For a more detailed explanation, see [27].
As mentioned, our goal is to obtain purely discrete versions of these approximations at grid nodes for our mimetic SBP-FD approximations. It is straight forward to use the relationships in section 2.1 to obtain analogous expressions for grad, curl, and div that are purely nodal-based. We do note however, that the underlying degrees of freedom are in general not only nodal, but also consist of line, area and volume integral degrees of freedom of the cell complex induced by the tensor product grid. The derivative matrix however gives us a way of computing nodal values from the integral degrees of freedom.
@@
$\frac{\partial{}}{\partial{x}}\phi(x_{i},y_{j},z_{k})=\sum_{s=1}^{N}V_{i,s}\left(\phi_{s,j,k}-\phi_{s-1,j,k}\right)=\sum_{s=0}^{N}D_{i,s}\phi_{s,j,k}.$ (24)
@@
In the following, we present discrete versions of all operators described by Gerritsma [15], which can be computed using arbitrary Lagrange and histopolation basis functions, for which we only need to know the function values and their derivatives at the grid nodes. On the other hand, as long as the values of the basis functions at the grid nodes are known, we can build the operators without having access to explicit basis functions. For instance, we can use high-order SBP-FD operators.
@@ 2.2.2 Curl Operator
We aim to define the curl operator in a space compatible with the gradient operator described above, so that we preserve the well-known identity that the curl of the gradient of a scalar field is zero. We choose $\tilde{\textbf{u}}=(\tilde{u}^{x},\tilde{u}^{y},\tilde{u}^{z})$ from the same space as $\textbf{grad}(\phi)$ defined above. That is, $\tilde{u}^{x}$ is from the tensor product space $U_{h}\otimes U_{l}\otimes U_{l}$, $\tilde{u}^{y}$ from $U_{l}\otimes U_{h}\otimes U_{l}$, and $\tilde{u}^{z}$ from $U_{l}\otimes U_{l}\otimes U_{h}$. Then the nodal values of $\textbf{curl}(\tilde{\textbf{u}})=\textbf{w}=(w^{x},w^{y},w^{z})^{T}$ are given by
$\begin{split}w^{x}(x_{i},y_{j},z_{k})&=\sum_{s=1}^{N}\sum_{t=1}^{N}\left(\tilde{u}_{i,s,t}^{z}-\tilde{u}_{i,s-1,t}^{z}-\tilde{u}_{i,s,t}^{y}+\tilde{u}_{i,s,t-1}^{y}\right)V_{j,s}V_{k,t}\\ &=\sum_{s=0}^{N}\sum_{t=1}^{N}\tilde{u}_{i,s,t}^{z}D_{j,s}V_{k,t}-\sum_{s=1}^{N}\sum_{t=0}^{N}\tilde{u}_{i,s,t}^{y}V_{j,s}D_{k,t},\end{split}$ (28)
$\begin{split}w^{y}(x_{i},y_{j},z_{k})&=\sum_{s=1}^{N}\sum_{t=1}^{N}\left(\tilde{u}_{s,j,t}^{x}-\tilde{u}_{s,j,t-1}^{x}-\tilde{u}_{s,j,t}^{z}+\tilde{u}_{s-1,j,t}^{z}\right)V_{i,s}V_{k,t}\\ &=\sum_{s=1}^{N}\sum_{t=0}^{N}\tilde{u}_{s,j,t}^{x}V_{i,s}D_{k,t}-\sum_{s=0}^{N}\sum_{t=1}^{N}\tilde{u}_{s,j,t}^{z}D_{i,s}V_{k,t},\end{split}$ (29)
$\begin{split}w^{z}(x_{i},y_{j},z_{k})&=\sum_{s=1}^{N}\sum_{t=1}^{N}\left(\tilde{u}_{s,t,k}^{y}-\tilde{u}_{s-1,t,k}^{y}-\tilde{u}_{s,t,k}^{x}+\tilde{u}_{s,t-1,k}^{x}\right)V_{i,s}V_{j,t}\\ &=\sum_{s=0}^{N}\sum_{t=1}^{N}\tilde{u}_{s,t,k}^{y}D_{i,s}V_{j,t}-\sum_{s=1}^{N}\sum_{t=0}^{N}\tilde{u}_{s,t,k}^{x}V_{i,s}D_{j,t}.\end{split}$ (30)
@@
$\begin{split}&u_{i,j,k}^{z}-u_{i,j-1,k}^{z}-u_{i,j,k}^{y}+u_{i,j,k-1}^{y}\\ =&\phi_{i,j,k}-\phi_{i,j,k-1}-\phi_{i,j-1,k}+\phi_{i,j-1,k-1}\\ -&\phi_{i,j,k}+\phi_{i,j-1,k}+\phi_{i,j,k-1}-\phi_{i,j-1,k-1}\\ =&0,\end{split}$ (31)